OpenStudy (anonymous):

Which of the following has the smallest value? 19% 0.9

OpenStudy (anonymous):

19%=19/100=0.19 so, 19%<0.9

OpenStudy (anonymous):

1/4 also and im thinking its 0.0.9

OpenStudy (anonymous):

0.9*

OpenStudy (anonymous):

If you convert .9 to a percentage, you would go two places to the right and realize it is 90%. Then you would see, 90% is a lot greater than 19% :)

OpenStudy (anonymous):

but i have 1/4 as well

OpenStudy (anonymous):

and thats 25

OpenStudy (anonymous):

so 19

OpenStudy (anonymous):

0.9 = 0.9*100%=90% now , 19% lesser than 90% so, 19% is smallest